OpenStudy (anonymous):

A sail is in the form of a right triangle that is four times as high as it is wide.The sail is made from 32 square meters of material.

OpenStudy (anonymous):

4 meters wide, 16 meters high

OpenStudy (anonymous):

What is the question?

OpenStudy (anonymous):

Does it ask for the height?

OpenStudy (anonymous):

Yes, It asks for the height.

OpenStudy (anonymous):

A is the area. H is the height, B is the base. Got it?

OpenStudy (anonymous):

A = 1/2 * B *H H=4B

OpenStudy (anonymous):

5x/2=32 5x=64 x=12.8 (width) 4x=51.2 (length)

OpenStudy (anonymous):

So, substitute 4B for H in the first equation to get A = 1/2 * B * 4B or A = 2 * B^2

OpenStudy (anonymous):

You following?

OpenStudy (anonymous):

Yeah, I got it. Thanks.

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

Ask if you have any more questions.